OpenStudy (anonymous):

You flip a coin three times. What is P(heads, heads, heads)? A. 18 B. 12 C. 13 D. 14

Mehek (mehek14):

are those fractions?

Mehek (mehek14):

or whole numbers?

OpenStudy (anonymous):

fractions sorry a) 1/8 b)1/2 c)1/3 d)1/4

Mehek (mehek14):

\(\dfrac{1}{2}*\dfrac{1}{2}*\dfrac{1}{2}=\dfrac{1}{8}\)

OpenStudy (anonymous):

oh okay thank you !

Mehek (mehek14):

np ^_^